# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v2

@@ 4.2 Pruning the Mixtral Architecture
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy [11] which just drops the experts less visited, router-guided strategy [21] has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies [26] which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers. We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
Results on Mixtral-8x22B. In this experiment, the model-centric method achieves the best result with only a minor performance drop of $2.8\%$ on average compared to the full model. Our proposed data-centric method achieves the second to the last performance, with a gap of $0.8\%$ performance gap compared with the runner-up method. It should be noted that the routing similarity-based pruning suffers model collapse under both two settings, showing a less robustness compared with other methods and metrics for similarity computation.
@@ 4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the Deepseek architecture. Specifically, we pruned the experts of Deepseek-MoE-16B, reducing the number of experts from 64 to 48 using different model pruning strategies. The results are illustrated in fig. 4.
@@ 4.5 Empirical Analysis on Expert Hints
Table 3: Evaluation results of different pruning setting with ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
@@ 4.6 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. We provide the full enumeration results in section A.1.
On the number of samples. We conduct experiments to study the effect of the used sample in data-centric pruning methods. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
We attribute this phenomenon to the use of mean operation in our expert similarity estimation. On the one hand, the mean operation keeps the feature dimension for efficient pairwise distance computation. On the other hand, it naturally downgrades the precision of function similarity computation by “round-off” error compared with other per sample-based distance metrics like CKA used in section 3.2.
On merging the routing policy. One relevant work to ours is [21], which identifies and merges similar experts based on routing policy, setting the routing policy of the merged expert to that of the most frequently visited expert. While we identify similar experts based on expert knowledge, such as weights and features, we merge not only the experts but also the routing policies. To demonstrate the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies. As shown in table 4, this results in performance degradation across all tasks and different numbers of experts, highlighting the crucial role of simultaneously merging the routers and experts.
Table 4: Evaluation results of our augmentation-based pruning method without merging the routers.
@@ 5 Conclusion and Future Work
In this paper, we address the practical problem of reducing the number of experts in MoE layers to accelerate the inference and fine-tuning process. We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible. This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
6 Acknowledgments
We would like to thank Lifeng Li from Microsoft for setting up GPU machines.
Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and fig. A2. We summarize the observation as follows,
@@ Appendix A Evaluation on the expert similarity
Figure A1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
Figure A2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
A.1 Enumeration on the expert pruning
We present the full greedy search result on Mixtral-8x7B in fig. A3. In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task. For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
@@ A.1 Enumeration on the expert pruning
Figure A3: Enumeration on dropping two experts.
A.2 Results on different strategies for pruning
@@
For model centric, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts. We uniformly merge the experts and their routers for model pruning. For data centric, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), or only maintain the expert with the maximum visiting frequency (Max).
From the results of our strategies, we can also notice two facts: 1) Both Vectorized $\theta$ and Surrogate $\theta$ strategies surpass all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning. 2) Compared with only using $\theta$, the involvement of $data$ in the pruning process can help better candidates for merging and pruning for improving the performance. Besides, though the learning strategy can bring a further improvement in the average performance, the computation cost is larger compared with using uniform coefficients for merging.
Table A1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Model-centric Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Surrogate $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
[…]
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Model-centric Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Surrogate $\bf{\theta}$ 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
[…]
Max $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7

### v3

@@ 4.2 Pruning the Mixtral Architecture
Results on Mixtral-8x7B. We can see that all our proposed four strategies surpass the related works, with a clear margin performance improvement of $1.5\%$ on average. Compared with count-guided strategy He et al. (2024) which just drops the experts less visited, router-guided strategy (Li et al., 2024a) has a large improvement of $3.7\%$ on average by merging these experts, showing that the merging operation plays a crucial role in preserving the expert knowledge. Besides, compared with count-guided and enumerate strategies (Lu et al., 2024) which all adopt the dropping strategy, we can see that directly leveraging the expert feedback rather than the routing frequency is more suitable for task-agnostic pruning in MoE layers. We can also notice that the model-centric method surpasses all the other data-involved pruning baseline methods. This suggests that weights already encode fruitful data information and can be deployed to group experts for pruning.
Results on Mixtral-8x22B. In this experiment, the model-centric approach achieves the best result, with only a minor performance drop of $2.8\%$ on average compared to the full model. Our proposed data-centric method ranks second to last, with a performance gap of $0.8\%$ compared to the runner-up method. This suggests that model-centric approaches exhibit better robustness when pruning models of different scales, while data-centric methods are more prone to overfitting on small calibration datasets (as evidenced by the collapse of the route-guided method in table 2).
@@ 4.3 Pruning the Deepseek Model
We also evaluate the effectiveness of our proposed pruning method on compressing the DeepSeek architecture. Unlike Mixtral-MoE, DeepSeek-MoE features a shared expert and incorporates more fine-grained experts at each MoE layer. Specifically, we pruned the non-shared experts of DeepSeek-MoE-16B, reducing the number of experts from 64 to 48 using various model pruning strategies. The results are illustrated in fig. 4***We only show the pruning results of the Deepseek and Qwen models on the most challenging MMLU task and the average performance across MMLU, BoolQ, OpenBookQA, and RTE. Full results can be found in supplementary..
@@ 4.5 Empirical Analysis on Expert Hints
Table 3: Evaluation results when ranging the number of samples for expert similarity estimation. The augmentation is disabled in this experiment.
@@ 4.6 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. We provide the full enumeration results in appendix C.
On the used samples. We conduct experiments to study the effect of both sample size and augmentation in data-centric pruning methods. As shown in table 3, increasing the number of samples from $128$ to $512$ without augmentation leads to a noticeable performance improvement of up to $0.6\%$. However, when the sample size is further increased to $2,048$, the average performance drops significantly to $61.6\%$. In contrast, when comparing this result with table 1, where $128$ samples were used with augmentation, we observe an average performance improvement of $0.3\%$. This demonstrates the importance of both sample size and augmentation in enhancing generalization during the pruning process, as discussed in section 3.4.
We attribute the performance drop with large sample sizes to the mean operation in our expert similarity estimation. While the mean preserves the feature dimension for efficient pairwise distance calculations, it reduces the precision of similarity estimates due to "round-off" errors, compared to per-sample metrics like CKA used in section 3.2.
On merging the routing policy.The motivation of our work is to diversify expert knowledge by merging similar experts. A key distinction of our approach is the merging of routing policies, which potentially directs more tokens to the resulting merged expert. To highlight the importance of merging routers, we evaluate the performance of our pruning method without merging the routing policies. As shown in table 4, compared with the results in table 1 and table 2,this results in performance degradation across all tasks and different numbers of experts, underscoring the crucial role of simultaneously merging both routers and experts.
Table 4: Evaluation results of our data-centric pruning method without merging the routers.
@@ 5 Conclusion and Future Work
In this paper, we work on the task-agnostic pruning of sparse MoEs. We propose discovering similar experts at the feature level and then merging them in the weight space for MoE pruning while preserving as much original expert knowledge as possible. This approach allows the MoE layer to maintain diverse experts with different knowledge, thereby efficiently reducing redundancy.
However, we didn’t account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
6 Limitations
Several unexplored questions remain in our project. First, we designed various strategies to prune the MoE, and we observed that different models require different strategies to achieve optimal post-pruning performance. It remains unclear what causes these performance differences across strategies. Second, while the learning strategy at the merging step can bring slightly performance improvement, the cost is also large. The question of how to efficiently find the optimal merging coefficients remains. Third, in our work, we prune the same experts across different MoE layers, despite each layer having varying levels of redundancy. A key question remains: how can we push MoE compression to its limits while maintaining acceptable performance?
Appendix A Experiment details
We conducted our experiments using four 80GB NVIDIA A100 GPUs. In our learning strategy, we used SGD as the optimizer to learn the coefficients for expert merging, with the learning rate set to $1\times 10^{-3}$. The coefficients were initialized as identity matrices and optimized over $50$ epochs. For baseline methods and models, all usage and distribution comply with the terms of their license, i.e., Mixtral (Apache License 2.0), Deepseek (MIT license), Qwen (Tongyi Qianwen license), lm evaluation harness (MIT license).
Appendix B Evaluation on the expert similarity
Table B1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Model-centric Vectorized 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
Surrogate 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
[…]
Uniform 56.2/47.8 72.7/57.2 52.1/42.7 68.1/54.0 85.6/83.2 32.8/28.6 68.6/66.5 62.3/54.3
Mixtral-8x22B 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Model-centric Vectorized 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
Surrogate 63.7/58.1 80.0/72.5 62.1/54.3 75.6/68.3 88.0/85.2 34.6/31.2 69.0/68.6 67.6/62.6
[…]
Max 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7
Uniform 62.3/57.8 78.5/69.7 60.2/51.3 73.4/64.2 87.6/83.1 35.8/33.2 71.1/68.1 67.0/61.1
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. D1 and fig. D2. We summarize the observation as follows,
@@ Appendix B Evaluation on the expert similarity
Appendix C Enumeration on the expert pruning
We present the full greedy search result on Mixtral-8x7B in fig. D3. In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task. For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
@@ Appendix C Enumeration on the expert pruning
Appendix D Results on different strategies for pruning
@@
  1. 1.
    For model-centric stratgies, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts. We uniformly merge the experts and their routers for model pruning. In other words, the model-centric strategies differ at the discovery stage.
  2. 2.
    For data-centric strategy, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), only maintain the expert with the maximum visiting frequency (Max), or uniformly merging different experts (Uniform). Thus, in this experiment, the data-centric strategies differ at the merging stage.
From the results of our strategies, we can observe the following:
  1. 1.
    Both the Vectorized and Surrogate strategies surpass all other data-involved pruning baseline methods. This suggests that weights already encode valuable data information, which can be utilized to group experts for pruning.
  2. 2.
    When pruning relatively small models (Mixtral-8x7B), the inclusion of data in the pruning process improves candidate selection for merging and pruning, leading to better performance compared to using only model weights. While the learning strategy offers a slight improvement in average performance, it comes at a higher computational cost compared to using uniform coefficients for merging.
  3. 3.
    For larger models (Mixtral-8x22B), the model-centric method outperforms the data-centric method. We argue that this is due to overfitting to the small calibration dataset during pruning, especially when dealing with a large number of parameters. The small calibration dataset cannot approximate the distribution of the pre-training dataset effectively. This is evident from the significant performance drop observed when using the Max strategy for pruning Mixtral-8x22B.
Figure D1: Evaluation of the expert similarity for different MoE layers in Mixtral-8x7B under the linear kernel-based CKA criteria.
Figure D2: Evaluation of the expert similarity for different MoE layers in Mixtral-8x22B under the linear kernel-based CKA criteria.
Figure D3: Enumeration on dropping two experts.
Appendix E Full results on pruning Deepseek and Qwen
We present the full results of pruning Deepseek-MoE-16 and Qwen2-57B-14A in fig. E4 and fig. E5. It is evident that our proposed data-centric and model-centric strategies outperform all baseline methods in most test cases. Additionally, when examining different evaluation tasks, the results on MMLU show a more reliable and consistent trend as the number of pruned experts increases. However, for smaller tasks such as the RTE dataset, we observe some randomness in the evaluation results due to the limited dataset size.We did not include the enumeration-based method in our comparison, as it is time-consuming and difficult to complete within a limited timeframe, especially when the number of experts is large in these models.
Figure E4: Results on pruning Deepseek-MoE-16B.
Figure E5: Results on pruning Qwen2-57B-14A.